# Instructing Large Language Models for Low-Resource Languages: A Systematic Study for Basque

### v1

@@ 8 Conclusions
This systematic study on instruction-tuning LLMs for Basque reveals several key strategies for low-resource language adaptation. We found that target language corpora are essential for effective learning. Employing bilingual (English and Basque) synthetic instructions yielded the most robust models whereas English-only instructions remain competitive. Crucially, starting from an instruction-tuned English model and adapting it to Basque proved more effective than training a base model for both language and instruction-following, or pre-training for language separately before instruction tuning. Our work contributes new Basque models, open instruction and human preference datasets, and methodological insights to guide future low-resource LLM development.
@@
Finally, this study primarily focuses on the initial instruction-tuning phase. While we did collect preference data, we did not extend our analysis to include preference alignment techniques. Including this additional phase would again have doubled the experimental workload and human evaluation requirements.
@@ Acknowledgements
This work has been partially supported by the Basque Government (Research group funding IT1570-22 and IKER-GAITU project), the Spanish Ministry for Digital Transformation and of Civil Service, and the EU-funded NextGenerationEU Recovery, Transformation and Resilience Plan (ILENIA project, 2022/TL22/00215335). The models were trained on the Leonardo supercomputer at CINECA under the EuroHPC Joint Undertaking, project EHPC-EXT-2024E01-042. Julen Etxaniz and Ekhi Azurmendi hold a PhD grant from the Basque Government (PRE_2024_2_0028 and PRE_2024_1_0035, respectively).
@@ A.1 English instructions
To generate the English synthetic instructions, we followed a technique called Magpie Xu et al. (2025). Briefly, it consists of letting the model generate text starting from the user’s prompt instead of the assistant’s response. See, for instance Fig. 5(a), where the model is asked to continue with the chat template immediately after the user header. We defined 5 different prompts to generate different kinds of instructions (see Fig. 5). We generated instructions sampling from the model using different 10 temperature values, ranging from 0.8 to 1.2. After generating the instructions, we apply the following filters:
@@
Remove duplicates: we remove those instructions that are exactly the same.
@@
Repetitive prompts or responses: we remove those instances that contain words that repeat more than 100 times.
@@
Poor quality instructions or responses: we remove those instances that were classified as poor quality by the model itself.
@@
Unfinished instructions: we noticed that some instructions ended with ":", meaning that the instructions was incomplete. We removed those as well.
As we experimented with models of two sizes and did not want to have any external influence in our experiments, we generated the instructions twice, once per model size. We later used the instructions generated from the 8B model to train the 8B models, and those from 70B to train the 70B model. The total size in tokens of the English instructions used for training is 350M tokens. We are releasing111111To be released upon acceptance. the generated instructions (filtered and unfiltered) for replicability purposes.
@@ A.2 Basque (translated) instructions
As discussed in Section 3.3, we generated our Basque instructions by translating the English instructions using Base EU. The prompt we used to perform the translation is shown in Fig. 6. As for the few-shot examples, we included 2 conversations of general instructions and 1 conversation for math, arithmetic, code and translation instruction types. These examples were sampled from the English instructions and post-edited manually. The exact examples are released in the GitHub repository.
@@ Appendix B Training details
colspec=X[3]X[2]X[2]X[2],
cells=l,font=,
row1=c,font=,
row2=l,font=,
column1=l,font=,
colsep=2pt,
rowsep=0pt
[…]
Base EN Base EU Instruct EN
Instructions EN
[…]
BAS EU I EN
[…]
Instructions EU
BAS EN I EU
BAS EU I EU
INS EN I EU
Instructions EN+EU
BAS EN I EN+EU
BAS EU I EN+EU
INS EN I EN+EU
Corpus EU
[…]
INS EN C EU
Corpus EU + Instructions EN
BAS EN C EU I EN
BAS EU C EU I EN
INS EN C EU I EN
Corpus EU + Instructions EU
BAS EN C EU I EU
BAS EU C EU I EU
INS EN C EU I EU
Corpus EU + Instructions EN+EU
BAS EN C EU I EN+EU
BAS EU C EU I EN+EU
INS EN C EU I EN+EU
[…]
colspec=lrrrr,
cells=font=,
row1=c,font=,
rowsep=0pt
[…]
Any backbone + I EU 64Gb A100 128 264.5h 94.26Kg
Any backbone + I EN+EU 64Gb A100 128 456.5h 162.69Kg
Any backbone + C EU 64Gb A100 128 1,730.1h 205.54Kg
Any backbone + C EU I EN 64Gb A100 128 1,932.8h 688.86Kg
Any backbone + C EU I EU 64Gb A100 128 2,016.0h 718.50Kg
Any backbone + C EU I EN+EU 64Gb A100 128 2,327.5h 829.53Kg
[…]
@@
Table 4 summarizes the training costs of our experiments (see Table 3 for the variants nomenclature). In total, we trained 19 models: one Base EU, seventeen Instruct EU variants, and one 70B Instruct EU. Due to unforeseen circumstances, the Base EU and Instruct EU models were trained using different frameworks and infrastructure. The Base EU model was trained with NeMo Harper et al. (2024) on 64GB H100 GPUs provided by MareNostrum 5. In contrast, all Instruct EU variants were trained using Fully Sharded Data-Parallel Zhao et al. (2023) on 64GB A100 GPUs provided by CINECA Leonardo. The total compute time across all experiments amounted to 39,879.1 GPU hours, which corresponds to 4,729.11 kg CO2eq, based on carbon intensity estimates from ElectricityMaps.121212At the time of the experiments: 0.297 kg/kWh for Italy and 0.157 kg/kWh otherwise, according to https://www.electricitymaps.com/.
colspec = lcc,
cells=font=,
row1 = font=,
rowsep=0pt
[…]
Total Batch Tokens 2M 4M

### v2

@@ 8 Conclusions
This systematic study on instruction-tuning LLMs for Basque reveals several key strategies for low-resource language adaptation. We found that target language corpora are essential for effective learning. Employing bilingual (English and Basque) synthetic instructions yielded the most robust models whereas English-only instructions remain competitive. Crucially, starting from an instruction-tuned English model and adapting it to Basque proved more effective than training a base model for both language and instruction-following, or pretraining for language separately before instruction tuning. Our work contributes new Basque models, open instruction and human preference datasets, and methodological insights to guide future low-resource LLM development.
@@
Third, this study primarily focuses on the initial instruction-tuning phase. While we did collect preference data, we did not extend our analysis to include preference alignment techniques. Including this additional phase would again have doubled the experimental workload and human evaluation requirements.
@@ Acknowledgements
This work has been partially supported by the Basque Government (Research group funding IT1570-22 and IKER-GAITU project), the Spanish Ministry for Digital Transformation and of Civil Service, and the EU-funded NextGenerationEU Recovery, Transformation and Resilience Plan (ILENIA project, 2022/TL22/00215335; and ALIA project). The models were trained on the Leonardo supercomputer at CINECA under the EuroHPC Joint Undertaking, project EHPC-EXT-2024E01-042. Julen Etxaniz and Ekhi Azurmendi hold a PhD grant from the Basque Government (PRE_2024_2_0028 and PRE_2024_1_0035, respectively).
@@ A.1 English instructions
To generate the English synthetic instructions, we followed the Magpie technique Xu et al. (2025). Briefly, it consists in letting the model generate text starting from the user’s prompt instead of the assistant’s response. See, for instance, Fig.˜6(a), where the model is asked to continue with the chat template immediately after the user header. We defined 5 prompts to generate different kinds of instructions (Fig.˜6), then sampled instructions from the model using 10 different temperature values ranging 0.8–1.2. After generating the instructions, we applied the following filters:
@@
Duplicates: keep unique instances.
@@
Repetitive prompts or responses: remove instances with a sequence of tokens repeated over 100 times.
@@
Poor quality prompts or responses: remove instances regarded as poor-quality by the model itself.
@@
Unfinished instructions: we noticed that some instructions ended with “:”, meaning that the instructions was incomplete. We removed those as well.
As we experimented with models of two sizes and did not want to have any external influence in our experiments, we generated the instructions twice, once per model size. We later used the instructions generated from the 8B model to train the 8B models, and those from 70B to train the 70B model. The total size in tokens of the English instructions used for training is 350M tokens. We are releasing101010hf.co/collections/HiTZ/latxa-instruct-682f356091452b0028380804 the generated instructions (filtered and unfiltered) for replicability purposes.
@@ A.2 Basque (translated) instructions
As discussed in Section˜3.3, we generated Basque instructions by translating the English instructions using Base EU. The prompt we used to perform the translation is shown in Fig.˜7. As for the few-shot examples, we included 2 conversations of general instructions and 1 conversation for each math, arithmetic, code and translation instruction types. These examples were sampled from the English instructions and post-edited manually. The exact examples are released in the GitHub repository.
@@ A.3 Examples
A.4 Manual evaluation
We conducted a manual analysis on a small sample of instructions to validate the translation pipeline. Two key insights emerged from this analysis:
The quality of the original English instructions is not always optimal, particularly for translation-related tasks.
The translated instructions are generally of high quality, often significantly better than NLLB translations for document-level, code-related, and math-related tasks. However, translations sometimes modify the intent slightly. For instance, the English instruction “Translate this sentence to English: […]” was translated into Basque as “Itzuli esaldi hau euskarara: […]”, which literally means “Translate this sentence to Basque: […]”. Nevertheless, the instruction remains well-adapted to the task as a whole and does not degrade overall quality.
In general terms, however, we concluded that reliably estimating instruction quality is non-trivial. While we are actively exploring this direction, it remains outside the current scope of the paper.
@@ Appendix B Training details
| Training Data | Backbone model | ||
| Base EN | Base EU | Instruct EN | |
| Instructions EN | Instruct EN | BAS EU I EN | Instruct EN |
| Instructions EU | BAS EN I EU | BAS EU I EU | INS EN I EU |
| Instructions EN+EU | BAS EN I EN+EU | BAS EU I EN+EU | INS EN I EN+EU |
| Corpus EU | Base EU | Base EU | INS EN C EU |
| Corpus EU + Instructions EN | BAS EN C EU I EN | BAS EU C EU I EN | INS EN C EU I EN |
| Corpus EU + Instructions EU | BAS EN C EU I EU | BAS EU C EU I EU | INS EN C EU I EU |
| Corpus EU + Instructions EN+EU | BAS EN C EU I EN+EU | BAS EU C EU I EN+EU | INS EN C EU I EN+EU |
| Model | Hardware | # GPU | GPU hours | CO2eq |
| Base EU | 64Gb H100 | 32 | 960.0h | 105.51Kg |
| Any backbone + I EN | 64Gb A100 | 128 | 192.0h | 22.81Kg |
| Any backbone + I EU | 64Gb A100 | 128 | 264.5h | 94.26Kg |
| Any backbone + I EN+EU | 64Gb A100 | 128 | 456.5h | 162.69Kg |
| Any backbone + C EU | 64Gb A100 | 128 | 1,730.1h | 205.54Kg |
| Any backbone + C EU I EN | 64Gb A100 | 128 | 1,932.8h | 688.86Kg |
| Any backbone + C EU I EU | 64Gb A100 | 128 | 2,016.0h | 718.50Kg |
| Any backbone + C EU I EN+EU | 64Gb A100 | 128 | 2,327.5h | 829.53Kg |
| INST EN,70B C EU I EN | 64Gb A100 | 256 | 16,005.1h | 1,901.41Kg |
| Total | - | - | 39,879.1h | 4,729.11Kg |
@@
Table˜4 summarizes the training costs of our experiments (see Table˜3 for the variants nomenclature). In total, we trained 19 models: one Base EU, seventeen Instruct EU variants, and one 70B Instruct EU. Due to unforeseen circumstances, the Base EU and Instruct EU models were trained using different frameworks and infrastructure. The Base EU model was trained with NeMo Harper et al. (2024) on 64GB H100 GPUs provided by MareNostrum 5. In contrast, all Instruct EU variants were trained using Fully Sharded Data-Parallel Zhao et al. (2023) on 64GB A100 GPUs provided by CINECA Leonardo. The total compute time across all experiments amounted to GPU hours, which corresponds to kg CO2eq, based on carbon intensity estimates from ElectricityMaps.111111At the time of the experiments: kg/kWh for Italy and kg/kWh otherwise, according to https://www.electricitymaps.com/.
| Hyperparameter | 8B Models | 70B Models |
| GPUs | 128 | 256 |
| Sequence Length | 8192 | |
| Gradient Accumulation | 1 | |
| Micro Batch Size | 2 | |
| Total Batch Tokens | 2M | 4M |
| Epochs | 4 | |
| Optimizer | AdamW | |
| 0.9, 0.95 | ||
| Scheduler | Cosine | |
| Cosine min LR ratio | 0.33 | |
| Learning rate | ||
| Warm-Up ratio | 0.1 | |
| Weight Decay | 0.1 | |
| Precision | BFloat16 | |
| FSDP Sharding Strategy | HYBRID | FULL |
| # Instructions | Epochs | Accuracy |
| 1M | 1 | 59.91 |
| 1M | 4 | 61.37 |
| 4M | 1 | 59.82 |
| 4M | 4 | 61.97 |
Hyperparameters.
Table˜5 outlines the key hyperparameters used during training. Both 8B and 70B model variants were trained with consistent configurations in terms of sequence length, batch size, optimizer settings, and learning rate schedules. The main differences lie in the number of GPUs and the FSDP sharding strategy, which was adjusted to better accommodate the increased memory and compute demands of the larger model. These hyperparameters were optimized using 8B model variant of Base EU, and the initial iteration on the instructed variants. We found this configuration to robustly perform across all the configurations.
We also explored the balance between the amount of English instructions and Basque monolingual data during joint training. We evaluated each setting on a subset of the development benchmarks, obtaining the results in Table˜6, which show that the number of training epochs had a more significant impact on performance than the number of instructions. However, to better assess instruction quality, we also conducted a small-scale internal arena evaluation. We observed that models trained with 4M instructions tended to produce worse responses. Based on these findings, we selected the 1M instructions + 4 epochs configuration as the most balanced setup. We have several hypotheses which would explain the results above:
1.
All examples are generated using the same method—Magpie—from the backbone model, and they tend to be quite homogeneous within their respective clusters general, translation, code, math, and arithmetic.
2.
Prior work by Etxaniz et al. (2024b) suggests that only a limited amount of English is needed during continual pretraining to maintain cross-lingual capabilities and avoid catastrophic forgetting. That is, 1M English instructions could be enough to avoid catastrophic forgetting when doing continual pretraining with Basque data.
Appendix C Evaluation details
C.1 Static Benchmarks
Our evaluation framework comprises a total of 27 benchmarks across three languages: 14 in Basque, 9 in English, and 4 in Spanish. These benchmarks span six categories designed to test different aspects of model capabilities:
Miscellaneous knowledge: BertaQA Etxaniz et al. (2024a), which tests knowledge of local Basque culture versus global topics; EusTrivia and EusExams from the Latxa suite Etxaniz et al. (2024b); and a subset of MMLU Hendrycks et al. (2021), manually translated to Basque Corral et al. (2025) and Spanish.121212hf.co/datasets/openai/MMMLU
Except MGSM, the datasets are framed as multiple-choice problems where models’ answers are determined by selecting the option with the highest log probability. MGSM is implemented as a generative task where an answer is directly sampled from the evaluated model and matched against a reference answer. We specifically chose the native chain-of-thought scenario. To provide models with contextual examples, our evaluations employed a 5-shot setting.
C.2 Arena Details
Guidelines for arena participants.
Fig.˜11 contains the information and instructions that were given to the human annotators who participated in the community-driven arena initiative. All the data collected through this initiative was properly anonymized prior to publication. Note that the actual information was provided in Basque, while here we show a translation to English.
Arena infrastructure.
On the infrastructure side, we used vLLM Kwon et al. (2023) to serve all model pairs and the baseline. We deployed a total of 18 endpoints for 8B models and one endpoint for a 70B model, running on nine and two A100 80GB GPUs, respectively. For the frontend, we developed a lightweight Gradio Abid et al. (2019) interface that allowed users to enter prompts, view model responses, and indicate their preferences based on content quality, language quality, or, in cases where no clear winner emerged, overall quality. To ensure a fair comparison across models, all models were given the same system prompt (shown in Fig.˜10) and the same hyperparameters: 0.9 temperature and 0.95 top-p.
Bradley–Terry model.
By using a scaling factor of 400, we ensure that the scores are interpretable in a manner consistent with the online ELO rating system. Thanks to its score stability and the assumption that model performance remains constant over time, the Bradley–Terry scoring system has become a widely adopted method for ranking LLMs—particularly since its introduction in the Chatbot Arena.131313lmarena.ai
C.3 Safety and Bias
Safety.
To test safety, we construct a Basque-language dataset inspired by XSTest Röttger et al. (2024), combining clearly unsafe prompts with superficially similar but safe ones. We measure both Violation rates (VR) and False Refusal rates (FRR) where the model wrongly declines safe prompts. The dataset includes a total of 50 instances, comprising both unsafe and safe prompts across five sensitive categories (self-harm, drugs, child-exploitation, terrorism, and explicit-content), adapted to the Basque context and translated into English for cross-language comparison. The outputs of the models were manually annotated by three members performing red teaming. Annotation agreement for unsafe prompt outputs was high (average agreement percentage: 0.973; Fleiss’ Kappa: 0.786; Krippendorff’s Alpha: 0.789). For safe prompt outputs, the annotators unanimously agreed on every item.
Bias.
For bias evaluation, we use BasqBBQ Zulaika and Saralegi (2025) for Basque and BBQ Parrish et al. (2022) for English to analyze disparities across languages. The evaluation is conducted in the same way as described in Section˜5.1, using LM Evaluation Harness framework with 4 few-shot examples. We use the accuracy metric to evaluate the bias of the model, measuring its ability to choose the correct answer even when biased traps are added to mislead it Parrish et al. (2022).
Appendix D Detailed Results
D.1 Benchmark Results
| ARC C | ARC E | Bele | BQA G | BQA L | BLMP | EusEx | EusPro | EusRe | EusTri | MGSM | MMLU | PIQA | XSC | Avg | |
| Base EN | 28.84 | 49.75 | 61.56 | 63.29 | 43.65 | 74.06 | 45.63 | 32.69 | 47.44 | 43.79 | 26.40 | 47.41 | 56.92 | 56.72 | 48.44 |
| + I EU | 36.35 | 61.95 | 69.44 | 67.39 | 42.98 | 84.06 | 48.39 | 35.42 | 44.89 | 45.42 | 30.80 | 50.37 | 61.44 | 62.14 | 52.93 |
| + I EN+EU | 38.28 | 63.50 | 72.00 | 68.76 | 42.77 | 84.25 | 48.41 | 36.19 | 45.17 | 46.27 | 37.20 | 52.04 | 62.09 | 62.84 | 54.27 |
| + C EU I EN | 39.42 | 64.81 | 73.00 | 70.99 | 63.49 | 91.28 | 51.92 | 48.66 | 56.82 | 56.33 | 47.20 | 51.85 | 67.32 | 69.36 | 60.89 |
| + C EU I EU | 38.14 | 67.00 | 71.78 | 71.24 | 62.65 | 92.39 | 49.88 | 47.20 | 46.02 | 57.78 | 39.20 | 50.00 | 65.63 | 66.71 | 58.97 |
| + C EU I EN+EU | 40.10 | 68.60 | 74.89 | 72.74 | 63.79 | 92.50 | 52.78 | 47.65 | 53.12 | 59.59 | 48.40 | 51.48 | 64.87 | 66.84 | 61.24 |
| Base EU | 38.65 | 67.38 | 75.22 | 72.45 | 65.65 | 92.50 | 55.03 | 53.57 | 58.24 | 60.52 | 36.00 | 51.11 | 65.09 | 68.50 | 61.42 |
| + I EN | 40.36 | 63.93 | 72.22 | 70.94 | 63.20 | 90.11 | 52.21 | 48.31 | 59.09 | 57.78 | 45.60 | 52.22 | 65.85 | 70.55 | 60.88 |
| + I EU | 39.25 | 67.09 | 73.56 | 71.78 | 61.59 | 92.11 | 52.38 | 47.78 | 49.43 | 57.14 | 35.60 | 53.70 | 64.92 | 67.31 | 59.55 |
| + I EN+EU | 38.31 | 66.92 | 74.56 | 72.58 | 62.23 | 91.22 | 53.04 | 48.96 | 53.98 | 58.95 | 44.80 | 56.30 | 64.27 | 67.70 | 60.99 |
| + C EU I EN | 40.96 | 66.12 | 61.22 | 71.49 | 63.83 | 91.72 | 50.67 | 47.19 | 47.16 | 57.49 | 43.60 | 52.59 | 66.83 | 70.35 | 59.37 |
| + C EU I EU | 37.97 | 67.42 | 71.33 | 71.24 | 62.52 | 91.83 | 51.76 | 46.31 | 52.84 | 57.32 | 37.60 | 52.22 | 65.90 | 67.31 | 59.54 |
| + C EU I EN+EU | 39.16 | 68.43 | 71.67 | 71.70 | 64.13 | 92.06 | 52.61 | 48.09 | 53.69 | 58.43 | 40.80 | 54.44 | 66.23 | 68.63 | 60.72 |
| Instruct EN | 29.10 | 50.88 | 73.89 | 67.10 | 44.97 | 69.61 | 46.21 | 34.13 | 49.72 | 45.01 | 45.60 | 50.37 | 57.63 | 61.22 | 51.82 |
| + I EU | 38.65 | 63.85 | 78.00 | 69.57 | 42.98 | 83.67 | 51.80 | 38.36 | 50.57 | 45.20 | 27.20 | 55.56 | 62.64 | 64.73 | 56.76 |
| + I EN+EU | 39.59 | 64.65 | 79.22 | 70.40 | 43.10 | 84.00 | 51.43 | 38.69 | 52.56 | 52.80 | 35.20 | 54.81 | 62.85 | 64.13 | 57.51 |
| + C EU | 37.97 | 65.70 | 77.33 | 73.87 | 66.33 | 92.67 | 55.05 | 52.12 | 58.24 | 61.40 | 48.40 | 51.85 | 66.99 | 70.28 | 62.73 |
| + C EU I EN | 41.38 | 66.79 | 80.00 | 74.62 | 65.23 | 91.39 | 56.00 | 52.83 | 59.66 | 61.05 | 54.00 | 57.04 | 67.32 | 71.34 | 64.19 |
| + C EU I EU | 40.44 | 69.15 | 81.44 | 73.54 | 66.07 | 91.83 | 56.01 | 52.06 | 62.78 | 62.33 | 46.40 | 52.96 | 66.01 | 71.01 | 63.72 |
| + C EU I EN+EU | 39.85 | 70.16 | 83.00 | 72.99 | 65.57 | 92.28 | 56.23 | 52.35 | 61.93 | 62.10 | 50.80 | 56.30 | 65.69 | 69.56 | 64.20 |
| Instruct EN,70B | 44.97 | 72.18 | 89.11 | 83.53 | 53.51 | 80.83 | 63.28 | 43.59 | 72.16 | 62.51 | 76.40 | 68.52 | 66.34 | 69.69 | 67.61 |
| + C EU I EN | 55.12 | 77.57 | 91.00 | 87.42 | 77.71 | 92.11 | 71.56 | 68.00 | 78.98 | 74.17 | 80.00 | 68.89 | 70.75 | 77.83 | 76.51 |
| 3.5 Sonnet | - | - | 94.22 | 93.52 | 80.46 | - | 82.68 | 81.60 | 87.39 | 84.61 | 85.20 | 79.63 | - | - | - |
| GPT-4o | - | - | 92.89 | 91.01 | 74.83 | - | 79.17 | 74.25 | 84.38 | 80.70 | 79.20 | 76.67 | - | - | - |
| English | Spanish | ||||||||||||||
| ARC C | ARC E | Bele | BQA G | BQA L | MGSM | MMLU | PIQA | XSC | Avg | Bele | MGSM | MMLU | XSC | Avg | |
| Base EN | 54.61 | 84.30 | 87.78 | 75.59 | 49.11 | 55.20 | 67.78 | 80.79 | 81.34 | 69.81 | 81.67 | 50.40 | 57.41 | 74.06 | 65.88 |
| + I EU | 54.95 | 83.63 | 86.78 | 73.95 | 47.38 | 67.60 | 63.70 | 80.63 | 82.46 | 70.73 | 80.67 | 58.40 | 57.78 | 74.12 | 67.74 |
| + I EN+EU | 54.82 | 81.42 | 87.61 | 74.46 | 47.91 | 78.00 | 61.30 | 80.25 | 82.40 | 72.52 | 82.00 | 67.60 | 58.89 | 75.28 | 70.94 |
| + C EU I EN | 53.84 | 80.72 | 85.89 | 75.75 | 60.41 | 79.20 | 60.00 | 79.71 | 82.40 | 73.03 | 81.33 | 63.60 | 60.00 | 73.53 | 69.62 |
| + C EU I EU | 54.44 | 82.20 | 80.33 | 75.08 | 59.56 | 62.00 | 55.19 | 80.69 | 80.21 | 69.12 | 75.22 | 58.00 | 53.70 | 72.01 | 64.73 |
| + C EU I EN+EU | 53.92 | 81.65 | 85.78 | 75.63 | 60.79 | 74.40 | 63.33 | 80.85 | 82.06 | 73.06 | 80.33 | 62.80 | 61.11 | 73.40 | 69.41 |
| Base EU | 52.39 | 82.83 | 84.33 | 76.05 | 61.72 | 55.20 | 63.70 | 80.30 | 80.08 | 68.87 | 79.11 | 46.80 | 58.52 | 72.20 | 64.16 |
| + I EN | 53.33 | 80.68 | 86.44 | 74.50 | 59.22 | 73.60 | 61.11 | 80.09 | 81.54 | 72.21 | 78.11 | 64.40 | 58.15 | 72.73 | 68.35 |
| + I EU | 54.44 | 82.41 | 84.11 | 74.87 | 58.63 | 64.00 | 62.22 | 80.09 | 81.27 | 70.65 | 77.22 | 56.40 | 54.81 | 74.19 | 65.66 |
| + I EN+EU | 53.75 | 80.89 | 85.89 | 74.41 | 59.14 | 73.20 | 63.70 | 79.22 | 81.14 | 72.44 | 80.44 | 64.40 | 56.67 | 74.39 | 68.97 |
| + C EU I EN | 54.18 | 81.94 | 84.11 | 74.50 | 59.48 | 74.40 | 59.63 | 80.14 | 81.27 | 71.65 | 76.33 | 56.80 | 57.41 | 73.33 | 65.97 |
| + C EU I EU | 53.92 | 82.53 | 81.78 | 74.50 | 59.09 | 58.00 | 62.96 | 80.09 | 80.21 | 69.09 | 74.89 | 47.60 | 55.56 | 72.60 | 62.66 |
| + C EU I EN+EU | 54.78 | 82.24 | 84.56 | 74.58 | 60.62 | 73.20 | 61.48 | 80.63 | 81.01 | 71.65 | 76.78 | 60.40 | 59.26 | 73.20 | 67.41 |
| Instruct EN | 57.76 | 85.48 | 92.67 | 77.47 | 50.51 | 87.20 | 66.67 | 81.28 | 83.52 | 77.02 | 87.89 | 78.80 | 62.96 | 77.50 | 76.79 |
| + I EU | 54.18 | 82.07 | 90.89 | 75.38 | 47.88 | 80.00 | 64.81 | 79.60 | 80.54 | 74.09 | 86.44 | 70.00 | 60.37 | 74.85 | 72.92 |
| + I EN+EU | 50.85 | 77.99 | 91.56 | 75.59 | 49.11 | 81.60 | 65.19 | 72.25 | 69.49 | 72.10 | 85.89 | 73.20 | 62.96 | 74.85 | 74.23 |
| + C EU | 53.84 | 81.86 | 90.22 | 77.05 | 63.16 | 68.40 | 64.81 | 80.20 | 83.26 | 73.74 | 82.67 | 61.20 | 59.63 | 73.20 | 69.17 |
| + C EU I EN | 51.96 | 79.42 | 91.00 | 77.38 | 62.01 | 81.20 | 62.59 | 74.59 | 76.31 | 74.35 | 85.67 | 73.20 | 59.26 | 72.47 | 72.65 |
| + C EU I EU | 53.58 | 81.65 | 90.67 | 76.55 | 61.63 | 72.00 | 64.07 | 79.60 | 80.28 | 74.65 | 84.22 | 70.00 | 58.52 | 75.18 | 71.98 |
| + C EU I EN+EU | 51.62 | 79.46 | 90.89 | 77.13 | 61.17 | 84.80 | 65.93 | 74.86 | 74.12 | 74.65 | 85.11 | 72.80 | 60.00 | 74.06 | 72.99 |
| Instruct EN,70B | 63.78 | 90.61 | 95.44 | 85.49 | 56.98 | 94.80 | 78.15 | 85.04 | 85.37 | 81.74 | 94.22 | 86.40 | 77.04 | 83.92 | 85.39 |
| + C EU I EN | 66.89 | 87.58 | 96.33 | 88.46 | 74.70 | 94.80 | 77.04 | 81.66 | 79.02 | 84.09 | 93.33 | 86.00 | 84.40 | 81.14 | 86.82 |
D.2 Arena Results
Complete human evaluation results in terms of Bradley–Terry scores and final rankings for each arena dimension (content, language, and global quality) can be consulted in Table˜9. Additionally, the detailed global win-rates and battle counts for each model pair are shown in Fig.˜12. Note that the reported win-rates incorporate ties in the calculation, resulting in win and loss percentages that do not necessarily sum to 100% for each model pair. Note also that base models (i.e., Base EN and Base EU) were not included in the arena evaluation, as they are not capable of following instructions. However, their performance on benchmarks can be consulted in Section˜D.1.
| Arena Global | Arena Content | Arena Language | |||||||||
| Rank | Bradley–Terry | Rank | Bradley–Terry | Rank | Bradley–Terry | ||||||
| GPT-4o | 1 | 1188 | +13/-17 | 1 | = | 1183 | +15/-13 | 1 | = | 1093 | +12/-10 |
| 3.5 Sonnet | 2 | 1153 | +13/-21 | 2 | = | 1150 | +12/-17 | 3 | 1 | 1082 | +11/-11 |
| INS EN,70B C EU I EN | 3 | 1141 | +15/-11 | 3 | = | 1127 | +10/-11 | 2 | 1 | 1083 | +13/-13 |